OpenStudy (anonymous):

i need help with dividing this problem: (x^4+x^3-10x^2+22x-8)/(x^2-3x+4)

OpenStudy (anonymous):

\[\frac{x^4+x^3-10x^2+22x-8}{x^2-3x+4}\] Just typesetting it.

OpenStudy (anonymous):

that is the correct problem...help...please :)

OpenStudy (anonymous):

You can do long division or you factor the numerator.

OpenStudy (anonymous):

could you show me how to factor the numerator? please

OpenStudy (anonymous):

\[(4 - 3 x + x^2) (-2 + 4 x + x^2) \]

OpenStudy (anonymous):

Your fraction becomes \[\frac{x^4+x^3-10 x^2+22 x-8}{x^2-3 x+4}=\frac {\left(x^2-3 x+4\right) \left(x^2+4 x-2\right) }{x^2-3 x+4} \] What is the answer?

OpenStudy (anonymous):

x^2+4x-2

OpenStudy (anonymous):

thanks...you made that look so easy....i was over analyzing....

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

will you chec to see if I got an answer correct?

OpenStudy (anonymous):

please

OpenStudy (anonymous):

(x^2+1/3x-2/9)/(x-1/3).....I got the answer of x-2/3 is that correct?

OpenStudy (anonymous):

i need help grahing a function: p(x)= -(x+2)(x-1)(x-3)